OpenStudy (anonymous):

an airplane flew for 5 hours with a tail wind of 25km/h. The return flight against the same wind took 5.5 hours. find the speed of the plane in still air

OpenStudy (anonymous):

D=RT THE DISTANCES ARE THE SAME THUS: RT=RT (X+25)5=(X-25)5.5 5X+125=5.5X-137.5 5X-5.5X=-137.5-125 -.5X=-262.5 X=-262.5/-.5 X=525 MPH IS THE PLANE SPEED IN STILL AIR.